OpenStudy (anonymous):

Let f(x) = 1/(3x)^1/2 . Compute lim h→0 (f(5 + h) − f(5))/h .

OpenStudy (anonymous):

I just need help simplfying the fraction after i plug in 5+h and 5

OpenStudy (anonymous):

After I plug it in the definition of derivative I get (1/(15+3h)^1/2)-(1/(15)^1/2))/h

zepdrix (zepdrix):

Ooo this one is a bit of a stinker >.<

zepdrix (zepdrix):

\[\Large\rm \lim_{h\to0}\frac{\frac{1}{\sqrt{15+3h}}-\frac{1}{\sqrt{15}}}{h}\]

OpenStudy (anonymous):

I tried combining the fractions and getting \[\frac{ \frac{ \sqrt{15}-\sqrt{15+3h} }{ \sqrt{15}\sqrt{15+3h} } }{ h }\]

OpenStudy (anonymous):

but because there are so many radicals, I'm having trouble taking this equation any further

OpenStudy (freckles):

\[\frac{1}{h} \frac{\sqrt{15}-\sqrt{15+3h}}{\sqrt{15}\sqrt{15+3h}} \cdot \frac{\sqrt{15}+\sqrt{15+3h}}{\sqrt{15}+\sqrt{15+3h}}\] try this...

OpenStudy (mathmath333):

\(\large\tt \color{black}{\lim_{h\to0}\dfrac{\dfrac{1}{\sqrt{15+3h}}-\dfrac{1}{\sqrt{15}}}{h}}\) \(\large\tt \color{black}{\lim_{h\to0}\dfrac{\dfrac{\sqrt{15}-(\sqrt{15+3h})}{(\sqrt{15})\sqrt{15+3h}}}{h}}\) \(\large\tt \color{black}{\lim_{h\to0}\dfrac{\dfrac{\sqrt{15}-(\sqrt{15+3h})}{(3)\sqrt{5(5+h)}}}{h}}\) \(\large\tt \color{black}{\lim_{h\to0}\dfrac{\dfrac{\sqrt{15}-(\sqrt{15+3h})}{(3)\sqrt{5(5+h)}}\times \dfrac{\sqrt{15}+(\sqrt{15+3h})}{(\sqrt{15}+(\sqrt{15+3h})}}{h}}\) \(\large\tt \color{black}{\lim_{h\to0}\dfrac{\dfrac{-h}{\sqrt{5(5+h)}(\sqrt{15}+(\sqrt{15+3h})}}{h}}\) \(\large\tt \color{black}{\lim_{h\to0}\dfrac{-1}{\sqrt{5(5+h)}(\sqrt{15}+(\sqrt{15+3h})}}\) \(\large\tt \color{black}{=\dfrac{-1}{\sqrt{5(5+0)}(\sqrt{15}+(\sqrt{15+3\times 0})}}\) \(\large\tt \color{black}{=\dfrac{-1}{10\sqrt{15}}}\)

OpenStudy (anonymous):

Oh wow! THANK YOU SO MUCH!!!!!!!!!

zepdrix (zepdrix):

Sorry had to go for a bit :c Looks like you've got it figured out though c: yay

OpenStudy (mathmath333):

y w